Make sure to show the steps, please. Thank you!2. Find the standard deviation of this probability distribution. Give your answer to at least 2 decimal places x P(x) 0 0.1 1
Answer #1

Solution:

x P(x) x * P(x) x2 * P(x)
0 0.1 0 0
1 0.25 0.25 0.25
2 0.3 0.6 1.2
3 0.35 1.05 3.15
Sum 1 1.9 4.6

Mean = \sum X * P(X)

\mu = 1.9

Standard deviation = \sigma

=\sqrt\sumX 2 * P(X) - \mu 2

=  \sqrt4.6 - (1.9)2 = 0.99

Standard deviation = \sigma = 0.99